OpenStudy (carol):

a car's gas tank can hold up to 20 gallons of gas. the tank currenlty has 16 gallons of gas. what percent of the tank is full?

OpenStudy (anonymous):

80%

OpenStudy (anonymous):

it is eighty percent because 16/20=4/5 which gives 400/5=80 percent. e_sanchez is right